A car with a mass 1,000kg is parked on a hill with a slope of 7.0°. The parking brake is released and the car starts to roll down the hill.
A) Assuming that friction is negligible, what will be the force that moves the car down the hill.
B) What will be the speed of the car after 50 seconds?
Answer #1

The force that moves the car down the hill is only the gravitational force that is along the slope

A)The fore that moves the car down is

F=mgsin(\theta)

=1000x9.81xsin(7)

=9810x0.12186

=1195.45 N

B) Velocity after 50 seconds is

V=at here a(acceleration)=gsin\theta

So V=(gsin(7degree))x50

=9.81x0.12186x50

=59.77 m/sec